# gather+[1][1]

### v1

@@ Abstract.
Matricized Tensor Times Khatri-Rao Product (MTTKRP) is the computational bottleneck in sparse tensor decomposition. As real-world sparse tensors grow to billions of nonzeros, they increasingly demand higher memory capacity and compute throughput from hardware accelerators. In this work, we present AMPED, a multi-GPU parallel algorithm designed to accelerate MTTKRP on billion-scale sparse tensors. AMPED scales beyond the limits of a single GPU, meeting both the memory and performance requirements of large-scale workloads. We introduce a partitioning strategy combined with a dynamic load balancing scheme to distribute computation and minimize GPU idle time. On real-world billion-scale tensors, AMPED achieves a 5.1$\times$ geometric mean speedup in total execution time over state-of-the-art GPU baselines using 4 GPUs on a single CPU node.
@@
copyright: nonejournalyear: 2025copyright: rightsretainedconference: 54th International Conference on Parallel Processing; September 8 - 11, 2025; San Diego, CAbooktitle: 54th International Conference on Parallel Processing (ICPP ’25), September 8 - 11, 2025, San Diego, CAccs: Computing methodologies Concurrent algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Vector / streaming algorithms
@@ 1. Introduction
  • We introduce a novel parallel algorithm to perform MTTKRP on sparse tensors using multiple GPUs. The proposed algorithm achieves geometric mean speedups of 1.9$\times$, 2.3$\times$, and 3.3$\times$ in execution time while using 2, 3, and 4 GPUs, compared to a single GPU implementation on billion-scale tensors.
  • We propose a tensor partitioning scheme to distribute the tensor elements across multiple GPUs. The proposed partitioning scheme achieves a geometric mean speedup of 8.2$\times$ in total execution time compared with equally distributing the nonzero tensor elements across the GPUs.
  • We introduce a load balancing scheme to distribute the computations of MTTKRP among GPUs while eliminating the communication of intermediate values across GPUs in each mode of computation. After each mode of computation, the updated factor matrix rows are shared across the GPUs. The proposed load balancing scheme shows less than $1\%$ overhead in computation time across GPUs on billion-scale tensors.
  • Our work achieves a geometric mean speedup of 5.1$\times$ in total execution time on billion-scale tensors compared with the state-of-the-art GPU baselines.
@@ 2.1. Introduction to Tensors
In this Section, we introduce terminology of tensors and tensor decomposition used in the paper. For a complete review of the tensors, see (Kolda and Bader, 2009).
A tensor is a multidimensional array whose dimensions are commonly referred to as modes. A $N$ mode tensor with real values is represented as $\mathcal{T}\in\mathbb{R}^{I_{0}\times\cdots I_{d}\times\cdots\times I_{N-1}}$, where the mode $d$ refers to the $d^{\text{th}}$ dimension of the tensor. Individual elements of the tensor, such as $x=\mathcal{T}(i_{0},i_{1},i_{2},...,i_{N-1})$, are indexed by $i_{0},i_{1},i_{2},...,i_{N-1}$, which are referred to as the indices of the nonzero tensor element. Tensors are often converted to a matrix form through a process known as matricization or matrix unfolding. The mode $d$ matricization of $\mathcal{T}$, denoted as $\mathcal{T}_{(d)}$, reorganizes the tensor into a matrix $\mathcal{T}_{(d)}\in\mathbb{R}^{I_{d}\times(I_{0}\cdots I_{d-1}I_{d+1}\cdots I_{N-1})}$. Here, the rows of the matricized tensor correspond to the indices of the mode $d$, while the columns are formed by iterating through all possible combinations of the indices in the remaining modes (Favier and de Almeida, 2014).
2.2. Canonical Polyadic Tensor Decomposition
In this Section, we illustrate CPD using a 3-mode tensor $\mathcal{T}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$ as an example.
CPD aims to approximate a tensor as a sum of rank-1 tensors. Specifically, given a tensor $\mathcal{T}$, the goal is to express it as:
$\mathcal{T}\approx\sum_{r=0}^{R-1}\mathbf{f{0}}_{r}\circ\mathbf{f{1}}_{r}\circ\mathbf{f{2}}_{r},$
where $R$ is a positive integer, and $\mathbf{f{0}}_{r}\in\mathbb{R}^{I_{0}}$, $\mathbf{f{1}}_{r}\in\mathbb{R}^{I_{1}}$, and $\mathbf{f{2}}_{r}\in\mathbb{R}^{I_{2}}$ for $r=0,\ldots,R-1$. Vectors $\mathbf{f{0}}_{r}$, $\mathbf{f{1}}_{r}$, and $\mathbf{f{2}}_{r}$ can be organized as factor matrices, $\mathbf{F_{0}}=[\mathbf{f{0}}_{0},\ldots,\mathbf{f{0}}_{R-1}]$, $\mathbf{F_{1}}$, and $\mathbf{F_{2}}$, respectively.
The rank of a tensor $\mathcal{T}$ is defined as the number of rank-1 tensors needed to exactly represent $\mathcal{T}$. When the rank is $R$, the CPD with $R$ components is computed by minimizing the Frobenius norm of the difference between the tensor and its approximation:
$\min_{\hat{\mathcal{T}}}\quad$ $\lVert\mathcal{T}-\hat{\mathcal{T}}\rVert^{2}_{F},$
subject to $\hat{\mathcal{T}}=[\![\lambda;\mathbf{F_{0}},\mathbf{F_{1}},\mathbf{F_{2}}]\!]=\sum_{r=0}^{R-1}\lambda_{r}\cdot\mathbf{f{0}}_{r}\circ\mathbf{f{1}}_{r}\circ\mathbf{f{2}}_{r},$
where $[\![\lambda;\mathbf{F_{0}},\mathbf{F_{1}},\mathbf{F_{2}}]\!]$ represents the Kruskal form of the tensor, and $\lambda_{r}$ is the weight of the $r$-th component.
The Alternating Least Squares (ALS) algorithm is the most commonly used method to compute the CPD due to its simplicity and computational efficiency. A critical kernel in CP-ALS is the Matricized Tensor-Times Khatri-Rao Product (MTTKRP), which dominates the computational cost. CP-ALS updates the factor matrices during each iteration by computing MTTKRP for each mode while keeping the other factor matrices fixed. For a 3-mode tensor, the MTTKRP operations are expressed as:
$\mathbf{\mathbf{F_{0}}}=\mathcal{T}_{(0)}(\mathbf{F_{1}}\odot\mathbf{F_{2}}),$
$\mathbf{\mathbf{F_{1}}}=\mathcal{T}_{(1)}(\mathbf{F_{2}}\odot\mathbf{F_{0}}),$
$\mathbf{\mathbf{F_{2}}}=\mathcal{T}_{(2)}(\mathbf{F_{0}}\odot\mathbf{F_{2}}),$
In this work, we focus on MTTKRP on sparse tensors, where the tensor $\mathcal{T}$ is sparse and the factor matrices $\mathbf{F_{0}}$, $\mathbf{F_{1}}$, and $\mathbf{F_{2}}$ are dense. Sparse MTTKRP is particularly challenging due to the irregular structure of the tensor, which demands efficient handling of sparsity to optimize performance.
@@ 3. Tensor Partitioning Scheme
3.1. Output Factor Matrix Row Update Computation (OFMRU)
Updating an output factor matrix row is the fundamental operation of MTTKRP. During MTTKRP along each mode (see Section 2.2), the output factor matrix rows are updated using each nonzero tensor element of the input tensor.
[…]
{gather+}
[0.85]  F_d(i_d, r) = F_d(i_d, r) + val(X(i_0, …, i_d, …, i_N-1)) ⋅∏_k = 0 k ≠d^N-1 F_k(i_k, r)
Following Section 2, Equation 3.1 shows the computation to update the output factor matrix, $\mathbf{F_{d}}$ using nonzero tensor element, $\mathcal{X}(i_{0},..,i_{d},..,i_{N-1})$ of a $N$ mode tensor. Here, $i_{0},..,i_{d},..,i_{N-1}$ is the position of the nonzero tensor element and $val(\mathcal{X}(i_{0},..,i_{d},..,i_{N-1}))$ is the value of the nonzero tensor element. Equation 3.1 computes the value in each column $r$ ($0\leq r<R$) in row $i_{d}$ of the mode $d$ factor matrix ($0\leq d<N$). Note that the row update is performed on all the nonzero tensor elements of the input tensor $\mathcal{X}$ to compute MTTKRP of output mode $d$.
@@ 3.2. Tensor Partitioning
We introduce a static partitioning scheme to partition an input tensor along each mode of an input tensor. Similarly to related works (Nisa et al., 2019c, a), we use multiple copies of the input tensor, where each tensor copy is partitioned targeting an output mode computation. Note that all tensor copies are stored in the large external memory of the host CPU. Our partitioning scheme in each mode distributes the nonzero tensor elements across (1) multiple GPUs using inter-device partitioning and (2) streaming multiprocessors (SMs) within each GPU using inter-SM partitioning. Note that an inter-device partition is executed by a single GPU, and an inter-SM partition is executed by a single GPU SM.
3.2.1. Inter-device Partitioning
The proposed partitioning scheme organizes the nonzero tensor elements of the input tensor into inter-device partitions (iDp) based on the output mode index along each mode. For a given output mode $d$, the OFMRU introduced in Section 3.1 (see Equation 3.1) is performed on each nonzero tensor element to update the corresponding rows of the output factor matrix. All nonzero tensor elements sharing the same output factor matrix index contribute to updating the same row. GPUs should maintain coherence among the updates that use nonzero tensor elements with the same output mode indices to avoid race conditions. To prevent race conditions between GPUs, all nonzero tensor elements sharing the same output mode index are assigned to the same $iDp$, eliminating task dependencies across GPUs. It avoids the need to maintain coherency across GPUs during execution, thereby significantly reducing synchronization overhead and inter-GPU data communication.
3.2.2. Inter-SM Partitioning
The nonzero tensor elements of each iDp are equally distributed among the SMs of each GPU by partitioning each iDp into equal sized inter-SM partitions (iSMp). With iSMp partitions, all the SMs of a GPU are assigned the same workload during execution. We use atomic operations to avoid race conditions between SMs of the same GPU.
@@ 3.3. Tensor Format Definition
Consider a multi-GPU setup with $m$ GPUs, each containing $g$ GPU SMs. For each output mode $d$, we divide the output mode indices $I_{d}$ into sets of equal sized partitions $I_{d,0},I_{d,1},\ldots,I_{d,k_{d}-1}$, where $k_{d}=\frac{|I_{d}|}{m}$. Here, $|I_{d}|$ denotes the size of $I_{d}$. Each index partition $I_{d,j}$ ($j=0,1,\ldots,(k_{d}-1)$) is a subset of the output mode indices $I_{d}$. Next, all nonzero tensor elements that incident on the indices in $I_{d,j}$ are collected into an inter-device partition, denoted by $\textit{{iDp}}_{d,j}$. To distribute nonzero tensor elements among GPU SMs, we further divide each inter-device partition into equal-sized sets called inter-SM partitions (iSMp). Each inter-device partition, $\textit{{iDp}}_{d,j}$ is divided into $t_{d,j}=\left\lceil|\textit{{iDp}}_{d,j}|/g\right\rceil$ inter-SM partitions. We denote the $q$-th inter-SM partition in $\textit{{iDp}}_{d,j}$ as $\textit{iSMp}_{d,j,q}$. For a tensor with $|\mathcal{T}|$ nonzero tensor elements, the total number of inter-SM partitions in mode $d$ is $\tau_{d}=\sum_{h=0}^{k_{d}-1}t_{d,h}\approx\frac{|\mathcal{T}|}{(m\times g)}$.
Observation 1.
Inter-device partitions (iDps) of a given output mode can be executed independently of other inter-device partitions of the same mode.
Proof.
Consider a nonzero tensor element $x_{i}\in\textit{iDp}_{d,j}$, where $\textit{iDp}_{d,j}$ is the $j^{\textrm{th}}$ inter-device partition of the input tensor $\mathcal{T}$ in mode $d$. Let the output mode index of $x_{i}$ in mode $d$ be $c_{d}$, meaning that $x_{i}$ contributes to updating the $c_{d}^{\textrm{th}}$ row of the output factor matrix of mode $d$ during the output mode $d$ MTTKRP computation.
From Equation 3.1 in Section 3.1, $x_{i}$ has data dependencies with any other tensor element $x_{z}$ if $x_{z}$ also updates the $c_{d}^{\textrm{th}}$ row of the output factor matrix. This implies that $x_{i}$ and $x_{z}$ share the same output mode index, $c_{d}$, and therefore contribute to the same output factor matrix row.
According to the tensor partitioning scheme defined in Section 3.2, all nonzero tensor elements with the same output mode index $c_{d}$ are assigned to the same iDp partition, $\textit{iDp}_{d,j}$. Thus, any $x_{z}$ with the same output mode index $c_{d}$ must also belong to $\textit{iDp}_{d,j}$ ($x_{z}\in\textit{iDp}_{d,j}$). Consequently, all nonzero tensor elements that update the same row of the output factor matrix are guaranteed to be in the same partition.
Since no two partitions share nonzero tensor elements with the same output mode index, there are no data dependencies between different iDp partitions of mode $d$. As a result, all $\textit{iDp}_{d,j}$ for a given mode $d$ can be executed independently. ∎
With Observation 1, iDps of each output mode can be executed on any GPU independently of each other.
@@ 4.2. Tensor Partition and Threadblock Mapping
Following the partition scheme proposed in Section 3.2, the inter-device partitions (iDp) and the inter-SM partitions (iSMp) are assigned to GPU Grids and threadblocks, respectively (Ansorge, 2022), as illustrated in Figure 1. Figure 1 uses the same notation introduced in Section 3.2. Once a GPU finishes executing all the computations in a Grid, a new Grid is loaded to the GPU for execution. Similarly, when a GPU SM finishes executing all the computations in a threadblock, a new threadblock in the same Grid is assigned to the SM.
@@ 4.4. Data Distribution Among GPUs
All the tensor partitions (of all output modes) are stored in the host CPU external memory. When an $iDp$ partition of the tensor is ready for execution, $iDp$ is transferred to the global memory of the corresponding GPU.
The factor matrices of each billion-scale tensor are significantly smaller (i.e., few mega-bytes) than the GPU global memory. Each GPU maintains a local copy of the factor matrices in its global memory. Once the MTTKRP computation for an output mode is completed, the updated rows of the output factor matrix are exchanged between GPUs to prepare for the computation of the following output mode.
@@ 4.5. Overall Algorithm
1 Input: Input tensor copies, $\textbf{T}=\{\mathcal{T}_{(0)},\mathcal{T}_{(1)}\cdots\mathcal{T}_{(N-1)}\}$
2 Factor matrices from previous iteration, $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$
3 Output: Updated factor matrices $\hat{\textbf{F}}$
4 for each mode $d=0,1,\ldots,(N-1)$ do
5       // $\mathcal{T}_{in}$ takes the partitions corresponding to $d^{\text{th}}$ matricization of the input tensor
6       $\mathcal{T}_{in}\leftarrow\mathcal{T}_{(d)}$

### v2

# AMPED: Accelerating MTTKRP for Billion-Scale Sparse Tensor Decomposition on Multiple GPUs

@@ Abstract.
[…]
@@
journalyear: 2025copyright: usgovmixedconference: 54th International Conference on Parallel Processing; September 08–11, 2025; San Diego, CA, USAbooktitle: 54th International Conference on Parallel Processing (ICPP ’25), September 08–11, 2025, San Diego, CA, USAdoi: 10.1145/3754598.3754651isbn: 979-8-4007-2074-1/25/09ccs: Computing methodologies Concurrent algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Vector / streaming algorithms
@@ 1. Introduction
[…]
@@ 2.1. Introduction to Tensors
Table 2. Notations
Symbol Details
$\circ$ vector outer product
$\otimes$ Kronecker product
$\odot$ Khatri-Rao product
$\mathbf{A}$ matrix
$\mathbf{a}$ vector
$a$ scalar
$\mathcal{X}$ sparse tensor
$\mathcal{X}_{(d)}$ mode-$d$ matricization of $\mathcal{X}$
A tensor is a generalization of an array to multiple dimensions. In the simplest high-dimensional case, a tensor is a three-dimensional array, which can be visualized as a data cube. For a thorough review of tensors, refer to (Kolda and Bader, 2009). Table 2 summarizes the tensor notations.
2.1.1. Tensor mode
In Tensor Decomposition, the number of dimensions of an input tensor is commonly called the number of tensor modes. For example, a vector can be seen as a mode-1 tensor. A $N$-mode, real-valued tensor is denoted by $\mathcal{X}\in\mathbb{R}^{I_{0}\times\cdots\times I_{N-1}}$. This paper focuses on tensors of mode three or higher for tensor decomposition.
2.1.2. Indices of a nonzero tensor element
For a 3-mode tensor, $\mathcal{X}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$, a nonzero tensor element is indicated as $x=\mathcal{X}(i_{0},i_{1},i_{2})$. Here, $i_{0}$, $i_{1}$, and $i_{2}$ are the positions or coordinates of $x$ in the tensor $\mathcal{X}$, which are commonly referred to as indices of the tensor element.
2.1.3. Tensor matricization
$\mathcal{X}_{(n)}$ denotes the mode-$n$ matricization or matrix unfolding (Favier and de Almeida, 2014) of $\mathcal{X}$. $\mathcal{X}_{(n)}$ is defined as the matrix $\mathcal{X}_{(n)}\in\mathbb{R}^{I_{n}\times(I_{0}\cdots I_{n-1}I_{n+1}\cdots I_{N-1})}$ where the parenthetical ordering indicates, the mode-$n$ column vectors are arranged by sweeping all the other mode indices through their ranges.
2.1.4. Canonical Poliyedic Tensor Decomposition (CPD)
CPD decomposes $\mathcal{X}$ into a sum of single-mode tensors (i.e., arrays), which best approximates $\mathcal{X}$. For example, given 3-mode tensor $\mathcal{X}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$, our goal is to approximate the original tensor as $\mathcal{X}\approx\sum_{r=0}^{R-1}\mathbf{a}_{r}\circ\mathbf{b}_{r}\circ\mathbf{c}_{r}$, where $R$ is a positive integer and $\mathbf{a}_{r}\in\mathbb{R}^{I_{0}}$, $\mathbf{b}_{r}\in\mathbb{R}^{I_{1}}$, and $\mathbf{c}_{r}\in\mathbb{R}^{I_{2}}$.
For each of the three modes, the spMTTKRP operation can be expressed as
[…]
The alternating least squares (ALS) method is used to compute CPD. In a 3-mode tensor, CPD sequentially performs the computations in Equation 1, iteratively. This can be generalized to higher mode tensors. Note that the matricization of $\mathcal{X}$ is different for each factor matrix computation. In this paper, performing MTTKRP on all the matricizations of an input tensor is called computing MTTKRP along all the modes. The outputs $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are the factor matrices that approximate $\mathcal{X}$. $\mathbf{a}_{r}$, $\mathbf{b}_{r}$, and $\mathbf{c}_{r}$ refers to the $r^{\text{th}}$ column of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$, respectively.
In this paper, we focus on MTTKRP on sparse tensors, which means that the tensor is sparse. Note that the factor matrices are dense.
@@ 3. Tensor Partitioning Scheme
3.0.1. Elementwise computation (EC)
Figure 1 summarizes the elementwise computation of a nonzero tensor element in mode 2 of a 3-mode tensor.
In Figure 1, the elementwise computation is carried out on a nonzero tensor element, denoted as $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$. In sparse tensors, $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ is typically represented in formats such as COOrdinate (COO) format. These formats store the indices ($i_{0}$, $i_{1}$, and $i_{2}$) along with the element value (i.e., $val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))$).
To perform the computation, $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ is first loaded onto the processing units (i.e., streaming multiprocessors for GPU) from the external memory (step 1). The compute device retrieves the rows $\mathbf{{A}}(i_{0},:)$, $\mathbf{{B}}(i_{1},:)$, and $\mathbf{{C}}(i_{2},:)$ from the factor matrices using the index values extracted from $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ (step 2, step 3, and step 4). Then, the compute device performs the following computation:
$\mathbf{{C}}(i_{2},r)=\mathbf{{C}}(i_{2},r)+val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))\cdot\mathbf{{A}}(i_{0},r)\cdot\mathbf{{B}}(i_{1},r)$
Here, $r$ refers to the column index of a factor matrix row ($r<R$). The operation involves performing a Hadamard product between row $\mathbf{{A}}(i_{0},:)$ and row $\mathbf{{B}}(i_{1},:)$, and then multiplying each element of the resulting product by $val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))$. Finally, the updated value is stored in the external memory (step 5).
Figure 1. Elementwise computation (Wijeratne et al., 2024)
@@ 3.1. Tensor Partitioning
We introduce a static partitioning scheme to partition an input tensor along each mode of the input tensor. Similarly to related works (Nisa et al., 2019c, a), we use multiple copies of the input tensor, where each tensor copy is partitioned targeting an output mode computation. Note that all tensor copies are stored in the large external memory of the host CPU.
Our partitioning scheme in each mode distributes the nonzero tensor elements across (1) multiple GPUs using tensor sharding and (2) streaming multiprocessors (SMs) within each GPU using inter-shard partitioning. Note that a tensor shard is executed by a single GPU, and an inter-shard partition is executed by a single GPU SM.
3.1.1. Tensor Sharding
The proposed sharding scheme organizes the nonzero tensor elements of the input tensor into tensor shards (TS) based on the output mode index along each mode. For a given output mode $d$, the EC introduced in Section 3.0.1 is performed on each nonzero tensor element to update the corresponding rows of the output factor matrix. All nonzero tensor elements sharing the same output factor matrix index contribute to updating the same row. GPUs should maintain coherence among the updates that use nonzero tensor elements with the same output mode indices to avoid race conditions. To prevent race conditions between GPUs, all nonzero tensor elements sharing the same output mode index are assigned to the same TS, eliminating task dependencies across GPUs. It avoids the need to maintain coherency across GPUs during execution, thereby significantly reducing synchronization overhead.
3.1.2. Inter-Shard Partitioning
The nonzero tensor elements of each tensor shard (TS) are equally distributed among the SMs of each GPU by partitioning each TS into equal-sized inter-shard partitions (ISP). With ISP partitions, all the SMs of a GPU are assigned the same workload during execution. We use atomic operations to avoid race conditions between SMs of the same GPU.
@@ 3.2. Tensor Format Definition
Consider a multi-GPU setup with $m$ GPUs, each containing $g$ GPU SMs. For each output mode $d$, we divide the output mode indices $I_{d}$ into sets of equal-sized partitions $I_{d,0},I_{d,1},\ldots,I_{d,k_{d}-1}$, where $k_{d}=\frac{|I_{d}|}{m}$. Here, $|I_{d}|$ denotes the size of $I_{d}$. Each index partition $I_{d,j}$ ($j=0,1,\ldots,(k_{d}-1)$) is a subset of the output mode indices $I_{d}$. Next, all nonzero tensor elements that incident on the indices in $I_{d,j}$ are collected into a tensor shard, denoted by $\textit{TS}_{d,j}$.
To distribute nonzero tensor elements among GPU SMs, we further divide each tensor shard into equal-sized sets called inter-shard partitions (ISP). Each tensor shard $\textit{TS}_{d,j}$ is divided into $t_{d,j}=\left\lceil|\textit{TS}_{d,j}|/g\right\rceil$ inter-shard partitions. We denote the $q$-th inter-shard partition in $\textit{TS}_{d,j}$ as $\textit{ISP}_{d,j,q}$. For a tensor with $|\mathcal{T}|$ nonzero tensor elements, the total number of inter-shard partitions in mode $d$ is $\tau_{d}=\sum_{h=0}^{k_{d}-1}t_{d,h}\approx\frac{|\mathcal{T}|}{(m\times g)}$.
@@ 4.2. Tensor Partition and Threadblock Mapping
Following the partition scheme proposed in Section 3.1, the tensor shards (TS) and the inter-shard partitions (ISP) are assigned to GPU Grids and threadblocks, respectively (Ansorge, 2022), as illustrated in Figure 2. Figure 2 uses the same notation introduced in Section 3.1. Once a GPU finishes executing all the computations in a Grid, a new Grid is loaded onto the GPU for execution. Similarly, when a GPU SM finishes executing all the computations in a threadblock, a new threadblock from the same Grid is assigned to the SM.
@@ 4.4. Data Distribution Among GPUs
All the tensor shards (for all output modes) are stored in the host CPU external memory. When a tensor shard ($TS$) is ready for execution, it is transferred to the global memory of the corresponding GPU.
The factor matrices of each billion-scale tensor are significantly smaller (i.e., a few megabytes) than the GPU global memory. Each GPU maintains a local copy of the factor matrices in its global memory. Once the MTTKRP computation for an output mode is completed, the updated rows of the output factor matrix are exchanged between GPUs to prepare for the computation of the next output mode.
@@ 4.5. Overall Algorithm
4.6. Mode-by-mode MTTKRP
Algorithm 1 shows the parallel algorithm for performing MTTKRP. Algorithm 1 takes (1) all the input tensor copies T and (2) factor matrices denoted as $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$. As shown in Algorithm 1, the MTTKRP is performed mode by mode (Algorithm 1: line 6). In each mode, A GPU grid (Algorithm 1: line 7) operates on a tensor shard mapped onto a GPU. At the end of all the computations of one mode, the GPUs are globally synchronized, and the generated output factor matrix rows are exchanged across GPUs before the computations of the next mode to maintain the correctness of the program (Algorithm 1: lines 8 - 11).
1 Input: Input tensor copies, $\textbf{T}=\{T_{0},T_{1}\cdots T_{N-1}\}$
2 Randomly initialized factor matrices, $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$
3 Output: Updated factor matrices $\hat{\textbf{Y}}=\{\hat{Y}_{0},\hat{Y}_{1},...,\hat{Y}_{N-1}\}$
4 for each mode $d=0,\ldots,N-1$ do
5 $T_{in}\leftarrow T_{d}$